OpenStudy (anonymous):

dy/dx=3x-2 where y=2 when x=-1

OpenStudy (nikita2):

y = 3/2*x^2 - 2*x + Const. Let's find the const: 2 = 3/2*(-1) - 2*(-1) + Const => Const = 3/2. Final: y(x) = 3/2x^2 - 2x + 3/2

OpenStudy (anonymous):

thak you